# From Perception to Reasoning: Deep Thinking Empowers Multimodal Large Language Models

### v1

@@ 3.1 CoT Paradigm
To activate the model’s deeper reasoning capabilities and simulate the logically rigorous ”System 2” thinking of humans, researchers have proposed the CoT mechanism and further extended the initial chain structure into various topological variants of reasoning paths(chu2023navigate), including typical forms such as chain structures(wei2022chain), tree structures(long2023large; yao2023tree), and graph structures(besta2024graph; lei2023boosting). These structures organize reasoning steps in different ways to improve the model’s capacity for multi-path thinking and reasoning flexibility. The core mechanism lies in decomposing a complex, one-step multimodal reasoning task into a series of simpler, more manageable intermediate steps, thereby reducing the risk of error accumulation in end-to-end generation.
@@
As shown in Figure 3, we compare the black-box reasoning approach of conventional MLLMs with representative textual and M-CoT paradigms. The chain-style CoT is the most fundamental reasoning structure, which generates intermediate steps in a linear sequence, where each reasoning result directly influences the next step. On this basis, Wang et al.(wang2022self) proposed CoT-SC, which introduces multiple independent reasoning chains. Due to the inherent uncertainty in large model reasoning, the same initial input may result in different outputs across reasoning chains. A predefined scoring function is then used to select the highest-scoring answer. The effectiveness of this strategy stems from its use of multi-path sampling and consistency checking, which effectively reduces ”compounding errors” caused by incidental mistakes in a single autoregressive path, thereby significantly enhancing the robustness of the results.
@@
In multimodal scenarios, most works still adopt a single chain of thought(yang2025r1; liu2024chain; gao2025interleaved), its primary value is in making the implicit cross-modal alignment process explicit. By generating intermediate language descriptions, it builds a bridge connecting visual perception and linguistic logic, thus effectively mitigating the representation gap between modalities. Some works have incorporated multimodal elements to further strengthen this bridge. For example, CoS(liu2024chain) method emphasizes the perception of key regions in an image to enhance visual reasoning, which essentially guides the model to anchor linguistic symbols to key visual evidence at each reasoning step. ICoT(gao2025interleaved) innovatively incorporates images as part of the reasoning content within intermediate steps. This represents a more thorough attempt to bridge the modality gap by creating a hybrid text-image semantic space where reasoning occurs directly within the multimodal information stream. Meanwhile, MVoT(li2025imagine) and Visual SKETCHPAD(hu2024visual) assist the model’s reasoning by generating visualizations of the process, which not only provides a visual scratchpad for the model but also offloads the cognitive load of purely linguistic symbolic reasoning through externalized visual representations.
@@
Other variants such as CoCoT(zhang2024cocot) enable MLLMs to compare similarities and differences across multiple image inputs and guide the model to answer detailed questions about them based on these findings. BBA(zhao2024bba) adopts a strategy similar to CoT-SC, constructing independent reasoning chains for the visual modality and domain-specific language (DSL) representations, and subsequently aligning and integrating them. Additionally, VISUALCODER (le2024visualcoder) combines CoT reasoning with visual control flow graphs for code reasoning tasks. To alleviate overthinking and structure the output, Xiang et al. proposed SCoT (xiang2025can), which adaptively generates various atomic steps while maintaining reasoning efficiency.
@@
For more complex modalities such as video and speech, VoT(fei2024video) was the first to successfully apply Chain-of-Thought techniques to achieve human-level video reasoning capabilities, demonstrating great potential across a wide range of video understanding scenarios. CoT-ST(du2024cot) significantly improved performance in speech translation tasks by introducing MCoT reasoning. The success of these methods on temporal modalities further highlights the advantage of task decomposition inherent in CoT. It breaks down continuous, high-dimensional video/speech signals into a series of manageable, text-based logical checkpoints, thereby making reasoning about complex dynamic scenes feasible.
@@
To systematically address this issue, the Tree of Thoughts (ToT) paradigm(long2023large; yao2023tree) was proposed, which mimics the human tree-like thought process for problem-solving and endows the model with the ability to backtrack. In a Tree of Thoughts, a single node represents a partial solution. For a given node, the model generates multiple child nodes (i.e., alternative options) and evaluates each one. Finally, search algorithms such as Breadth-First Search (BFS) or Depth-First Search (DFS) are used to determine the expansion of the tree. The essence of ToT is to expand a single, fragile reasoning path into a robust and explorable solution space. Its evaluation and search mechanisms are equivalent to introducing active pruning and error correction capabilities, thereby systematically overcoming the compounding error problem of the linear chain. Additionally, the BoT framework(chen2024boosting) iteratively explores and evaluates the tree of thoughts to gain experience from trial and error, thus arriving at a precise solution. The ToT paradigm allows the chain of thought to extensively explore the problem’s solution space and achieve local or even global optimal solutions.
The Graph of Thoughts (GoT)(besta2024graph; lei2023boosting) introduces a more complex topological structure, modeling a problem’s concepts as graph nodes and using edges to represent the relationships between them. Unlike a tree of thoughts, a graph of thoughts allows nodes to have multiple parent nodes, introducing cycles and N-to-1 connections. This enables the model not only to explore different reasoning paths but also to aggregate and synergize multiple parallel reasoning results, thus forming a more comprehensive solution. In a graph of thoughts, each subgraph represents the solving process of a subproblem, and the combination of all subgraphs constitutes the final solution. If ToT explores in depth and breadth, then GoT adds the capability to merge, constructing a more flexible and powerful non-linear reasoning structure. In multimodal scenarios, BDoG(zheng2024picture) constrains the debate process within a graph it refers to as a ”blueprint,” storing viewpoints and evidence in graph branches to reduce interference from frequent yet irrelevant concepts. Furthermore, Hypergraph-of-Thought (HoT)(yao2023thinking) models higher-order relationships by introducing hyperedges, further enhancing the ability to align and integrate complex cross-modal semantic relationships.
@@ 3.2 Post-Training Phase
COCO-MMRwei2024enhancing 2023 VQA Multi-domain T, I Open 62,351
$M$-PhyQAanand2024mm 2024 Physics QA Physics T, I Open 4,500
HaloQuestwang2024haloquest 2024 VQA Hallucination T, I Open 7.7K
MathV360Kshi2024math 2024 Math QA Math T, I Open 360K
CMM-Mathliu2024cmm 2024 Math QA Math T, I MC, Completion, Open 28K
Visual CoTshao2024visual 2024 VQA Multi-domain T, I Open 438K
GPSM4Kanand2024improving 2024 Math QA Geometry T, I Open 4,440
AMATH-SFTxiang2024atomthink 2024 Math QA Math T, I MC, Open 157K
LLaVA-CoT-100kxu2024llava 2024 VQA Commonsense, Science T, I MC, Open 834K
MLRQAxiao2024mlrqa 2025 VQA Multi-domain T, I MC 4,356
$M$-Verifysun2025mm 2025 Math QA Math T, I MC, Open 59,772
MMathCoT1Mluo2025ursa 2025 Math QA Math T, I MC 1M
CoTAxie2025audio 2025 Audio QA Multi-domain T, A Open 1.2M
Vision-R1-coldhuang2025vision 2025 VQA Multi-domain T, I MC, Open 200K
VisualPRM400Kwang2025visualprm 2025 Science QA Math, Science T, I MC, Open 400K
R1-Onevisionyang2025r1 2025 Multi-domain Multi-domain T, I MC, Open 155K
LongPerceptual Thoughtsliao2025longperceptualthoughts 2025 VQA Visual Perception T, I MC 30K
DriveLMM-o1ishaq2025drivelmm 2025 VQA Autonomous Driving T, I, PC MC, Open 18K
Emma-Xsun2024emma 2024 Robot Manipulation Embodied AI T, V Robot Action 60K
Embodied-Reasonerzhang2025embodied 2025 Robot Manipulation Embodied AI T, I MC, Open 9,390
@@ 3.2.1 Training Data
The COCO-MMR datasetwei2024enhancing extracts a large number of open-ended questions, reasoning chains, and answers from the COCO datasetlin2014microsoft, and pioneers the use of open-ended questions in the context of multimodal Chain-of-Thought reasoning to introduce new challenges for training and evaluating reasoning capabilities. MM-PhyQAanand2024mm comprises well-structured high-school-level multimodal physics problems and enhances model performance on multi-step physical reasoning tasks. HaloQuestwang2024haloquest focuses on capturing various aspects of multimodal hallucinations, contributing over 7.7K examples to help mitigate hallucination issues in multimodal reasoning models. Datasets such as MathV360Kshi2024math, AMATH-SFTxiang2024atomthink, and MMathCoT1Mluo2025ursa provide high-quality data for multimodal mathematical reasoning. Additionally, CMM-Mathliu2024cmm offers a Chinese-language multimodal math reasoning dataset, while GPSM4Kanand2024improving delivers detailed step-by-step solutions in a unified format for multimodal geometric reasoning tasks.
Datasets including Visual CoTshao2024visual, LLaVA-CoT-100Kxu2024llava, Vision-R1-coldhuang2025vision, and R1-Onevisionyang2025r1 support visual reasoning tasks across various domains and contain detailed multi-step reasoning annotations. MLRQAxiao2024mlrqa, composed by domain experts, includes logic reasoning data designed to activate the reasoning potential of MLLMs using multimodal Chain-of-Thoughts. Sun et al.sun2025mm propose two effective methods for synthesizing multimodal CoT datasets, offering practical solutions for generating large-scale multimodal reasoning data. For audio data, CoTAxie2025audio provides multi-task audio datasets with structured Chain-of-Thought reasoning processes.
@@
Given that most current CoT datasets focus on mathematically rigorous and code-based reasoning tasks, some researchers have begun exploring CoT data for other application scenarios. This expansion is of crucial importance because it demonstrates that CoT is not merely a specialized technique applicable only to mathematical logic, but rather a general framework with wide applicability that simulates the complex cognitive processes of humans. LongPerceptualThoughtsliao2025longperceptualthoughts aims to fill the gap in reasoning datasets for visual perception tasks, providing a dataset of 30K perceptual reasoning samples with long-form CoT traces. Meanwhile, Emma-Xsun2024emma and Embodied-Reasonerzhang2025embodied supply training data for embodied intelligence models. Emma-X offers a hierarchical embodied dataset containing 60K robot operation trajectories, while Embodied-Reasoner provides 9.3K coherent observation–thought–action trajectories. DriveLMM-o1 ishaq2025drivelmm contributes over 18K visual question-answering instances for step-by-step visual reasoning in autonomous driving, covering tasks such as perception, prediction, and planning.
@@ 3.2.2 Training Paradigm
To address this limitation and enhance learning efficiency, the researchers incorporated the scaffolding theory from cognitive science and guided the model to gradually acquire the reasoning ability. For instance, LlamaV-o1thawakar2025llamav adopts a multi-step curriculum learning strategy where the model first learns basic subtasks before progressing to multi-step reasoning training. Similarly, CoR(yu2025chain) and LLaVA-CoT(xu2024llava) employ a Progressive Paradigm Training (PPT) approach, introducing different types of reasoning data in stages to gradually expand the model’s reasoning capabilities.
@@
Unlike conventional MLLMs that use Reinforcement Learning from Human Feedback (RLHF) to align with human preferences and reduce hallucinations, reinforcement learning in CoT-MLLMs focuses more on improving the logicality, completeness, and problem-solving effectiveness of reasoning chains. This training paradigm typically employs reinforcement fine-tuning (RFT) methods based on Direct Preference Optimization (DPO)(rafailov2023direct) or Group Relative Policy Optimization (GRPO)(shao2024deepseekmath).
For example, Seg-Zero(liu2025seg) uses five manually designed reward functions to enable the model to acquire a reasonable level of reasoning ability without relying on any supervised reasoning data. Inspired by the cold-start strategy of DeepSeek-R1(guo2025deepseek), Vision-R1(huang2025vision) proposes a Progressive Thinking Suppression Training (PTST) strategy to alleviate the optimization challenge of excessive reasoning in early-stage reinforcement learning. This method restricts reasoning length initially and gradually relaxes this constraint during training, enabling the model to autonomously learn to use longer reasoning chains to solve complex problems, thereby enhancing its reasoning capabilities. In addition, Visual-RFT(liu2025visual) demonstrates that reinforcement fine-tuning can be effectively applied beyond mathematics and code to visual perception tasks, and it is the first to adopt Reinforcement Learning with Verifiable Rewards (RLVR) in image-based multimodal models. R1-Omni(zhao2025r1) further extends RLVR to video-based multimodal models for the first time.
@@
ARES(byun2024ares) proposes a two-stage algorithm that alternates between reinforcement learning and supervised fine-tuning. First, a teacher model scores the outputs to guide reinforcement learning; then, the teacher corrects erroneous reasoning steps, followed by supervised fine-tuning to stabilize post-RL performance. Moreover, Insight-V(dong2025insight) adopts an iterative DPO strategy in the reinforcement learning phase. By conducting multiple rounds of DPO training and sampling, the model better simulates the online DPO process, effectively addressing the issue of preference data drifting from the model distribution in offline generation. To further enhance the stability of multimodal reasoning models, Wang et al.(wang2024enhancing) introduce a Mixed Preference Optimization (MPO) approach, whose loss function consists of three components: preference loss, quality loss, and generation loss.

### v2

@@ 3.1 CoT Paradigm
To activate the model’s deeper reasoning capabilities and simulate the logically rigorous ”System 2” thinking of humans, researchers have proposed the CoT mechanism and further extended the initial chain structure into various topological variants of reasoning paths(Chu et al., 2023), including typical forms such as chain structures(Wei et al., 2022), tree structures(Long, 2023; Yao et al., 2023b), and graph structures(Besta et al., 2024; Lei et al., 2023). These structures organize reasoning steps in different ways to improve the model’s capacity for multi-path thinking and reasoning flexibility. The core mechanism lies in decomposing a complex, one-step multimodal reasoning task into a series of simpler, more manageable intermediate steps, thereby reducing the risk of error accumulation in end-to-end generation.
@@
As shown in Figure 3, we compare the black-box reasoning approach of conventional MLLMs with representative textual and M-CoT paradigms. The chain-style CoT is the most fundamental reasoning structure, which generates intermediate steps in a linear sequence, where each reasoning result directly influences the next step. On this basis, Wang et al.(Wang et al., 2022) proposed CoT-SC, which introduces multiple independent reasoning chains. Due to the inherent uncertainty in large model reasoning, the same initial input may result in different outputs across reasoning chains. A predefined scoring function is then used to select the highest-scoring answer. The effectiveness of this strategy stems from its use of multi-path sampling and consistency checking, which effectively reduces ”compounding errors” caused by incidental mistakes in a single autoregressive path, thereby significantly enhancing the robustness of the results.
@@
In multimodal scenarios, most works still adopt a single chain of thought(Yang et al., 2025c; Liu et al., 2024c; Gao et al., 2025), its primary value is in making the implicit cross-modal alignment process explicit. By generating intermediate language descriptions, it builds a bridge connecting visual perception and linguistic logic, thus effectively mitigating the representation gap between modalities. Some works have incorporated multimodal elements to further strengthen this bridge. For example, CoS(Liu et al., 2024c) method emphasizes the perception of key regions in an image to enhance visual reasoning, which essentially guides the model to anchor linguistic symbols to key visual evidence at each reasoning step. ICoT(Gao et al., 2025) innovatively incorporates images as part of the reasoning content within intermediate steps. This represents a more thorough attempt to bridge the modality gap by creating a hybrid text-image semantic space where reasoning occurs directly within the multimodal information stream. Meanwhile, MVoT(Li et al., 2025a) and Visual SKETCHPAD(Hu et al., 2024b) assist the model’s reasoning by generating visualizations of the process, which not only provides a visual scratchpad for the model but also offloads the cognitive load of purely linguistic symbolic reasoning through externalized visual representations.
@@
Other variants such as CoCoT(Zhang et al., 2024a) enable MLLMs to compare similarities and differences across multiple image inputs and guide the model to answer detailed questions about them based on these findings. BBA(Zhao et al., 2024a) adopts a strategy similar to CoT-SC, constructing independent reasoning chains for the visual modality and domain-specific language (DSL) representations, and subsequently aligning and integrating them. Additionally, VISUALCODER (Le et al., 2024) combines CoT reasoning with visual control flow graphs for code reasoning tasks. To alleviate overthinking and structure the output, Xiang et al. proposed SCoT (Xiang et al., 2025), which adaptively generates various atomic steps while maintaining reasoning efficiency.
@@
For more complex modalities such as video and speech, VoT(Fei et al., 2024) was the first to successfully apply Chain-of-Thought techniques to achieve human-level video reasoning capabilities, demonstrating great potential across a wide range of video understanding scenarios. CoT-ST(Du et al., 2024) significantly improved performance in speech translation tasks by introducing MCoT reasoning. The success of these methods on temporal modalities further highlights the advantage of task decomposition inherent in CoT. It breaks down continuous, high-dimensional video/speech signals into a series of manageable, text-based logical checkpoints, thereby making reasoning about complex dynamic scenes feasible.
@@
To systematically address this issue, the Tree of Thoughts (ToT) paradigm(Long, 2023; Yao et al., 2023b) was proposed, which mimics the human tree-like thought process for problem-solving and endows the model with the ability to backtrack. In a Tree of Thoughts, a single node represents a partial solution. For a given node, the model generates multiple child nodes (i.e., alternative options) and evaluates each one. Finally, search algorithms such as Breadth-First Search (BFS) or Depth-First Search (DFS) are used to determine the expansion of the tree. The essence of ToT is to expand a single, fragile reasoning path into a robust and explorable solution space. Its evaluation and search mechanisms are equivalent to introducing active pruning and error correction capabilities, thereby systematically overcoming the compounding error problem of the linear chain. Additionally, the BoT framework(Chen et al., 2024d) iteratively explores and evaluates the tree of thoughts to gain experience from trial and error, thus arriving at a precise solution. The ToT paradigm allows the chain of thought to extensively explore the problem’s solution space and achieve local or even global optimal solutions.
The Graph of Thoughts (GoT)(Besta et al., 2024; Lei et al., 2023) introduces a more complex topological structure, modeling a problem’s concepts as graph nodes and using edges to represent the relationships between them. Unlike a tree of thoughts, a graph of thoughts allows nodes to have multiple parent nodes, introducing cycles and N-to-1 connections. This enables the model not only to explore different reasoning paths but also to aggregate and synergize multiple parallel reasoning results, thus forming a more comprehensive solution. In a graph of thoughts, each subgraph represents the solving process of a subproblem, and the combination of all subgraphs constitutes the final solution. If ToT explores in depth and breadth, then GoT adds the capability to merge, constructing a more flexible and powerful non-linear reasoning structure. In multimodal scenarios, BDoG(Zheng et al., 2024a) constrains the debate process within a graph it refers to as a ”blueprint,” storing viewpoints and evidence in graph branches to reduce interference from frequent yet irrelevant concepts. Furthermore, Hypergraph-of-Thought (HoT)(Yao et al., 2023a) models higher-order relationships by introducing hyperedges, further enhancing the ability to align and integrate complex cross-modal semantic relationships.
@@ 3.2 Post-Training Phase
COCO-MMRWei et al. (2024) 2023 VQA Multi-domain T, I Open 62,351
MM-PhyQAAnand et al. (2024b) 2024 Physics QA Physics T, I Open 4,500
HaloQuestWang et al. (2024j) 2024 VQA Hallucination T, I Open 7.7K
MathV360KShi et al. (2024) 2024 Math QA Math T, I Open 360K
CMM-MathLiu et al. (2024b) 2024 Math QA Math T, I MC, Completion, Open 28K
Visual CoTShao et al. (2024a) 2024 VQA Multi-domain T, I Open 438K
GPSM4KAnand et al. (2024a) 2024 Math QA Geometry T, I Open 4,440
AMATH-SFTXiang et al. (2024a) 2024 Math QA Math T, I MC, Open 157K
LLaVA-CoT-100kXu et al. (2024a) 2024 VQA Commonsense, Science T, I MC, Open 834K
MLRQAXiao et al. (2024a) 2025 VQA Multi-domain T, I MC 4,356
MM-VerifySun et al. (2025c) 2025 Math QA Math T, I MC, Open 59,772
MMathCoT1MLuo et al. (2025) 2025 Math QA Math T, I MC 1M
CoTAXie et al. (2025) 2025 Audio QA Multi-domain T, A Open 1.2M
Vision-R1-coldHuang et al. (2025) 2025 VQA Multi-domain T, I MC, Open 200K
VisualPRM400KWang et al. (2025d) 2025 Science QA Math, Science T, I MC, Open 400K
R1-OnevisionYang et al. (2025c) 2025 Multi-domain Multi-domain T, I MC, Open 155K
LongPerceptual ThoughtsLiao et al. (2025) 2025 VQA Visual Perception T, I MC 30K
DriveLMM-o1Ishaq et al. (2025) 2025 VQA Autonomous Driving T, I, PC MC, Open 18K
Emma-XSun et al. (2024c) 2024 Robot Manipulation Embodied AI T, V Robot Action 60K
Embodied-ReasonerZhang et al. (2025e) 2025 Robot Manipulation Embodied AI T, I MC, Open 9,390
@@ 3.2.1 Training Data
The COCO-MMR datasetWei et al. (2024) extracts a large number of open-ended questions, reasoning chains, and answers from the COCO datasetLin et al. (2014), and pioneers the use of open-ended questions in the context of multimodal Chain-of-Thought reasoning to introduce new challenges for training and evaluating reasoning capabilities. MM-PhyQAAnand et al. (2024b) comprises well-structured high-school-level multimodal physics problems and enhances model performance on multi-step physical reasoning tasks. HaloQuestWang et al. (2024j) focuses on capturing various aspects of multimodal hallucinations, contributing over 7.7K examples to help mitigate hallucination issues in multimodal reasoning models. Datasets such as MathV360KShi et al. (2024), AMATH-SFTXiang et al. (2024a), and MMathCoT1MLuo et al. (2025) provide high-quality data for multimodal mathematical reasoning. Additionally, CMM-MathLiu et al. (2024b) offers a Chinese-language multimodal math reasoning dataset, while GPSM4KAnand et al. (2024a) delivers detailed step-by-step solutions in a unified format for multimodal geometric reasoning tasks.
Datasets including Visual CoTShao et al. (2024a), LLaVA-CoT-100KXu et al. (2024a), Vision-R1-coldHuang et al. (2025), and R1-OnevisionYang et al. (2025c) support visual reasoning tasks across various domains and contain detailed multi-step reasoning annotations. MLRQAXiao et al. (2024a), composed by domain experts, includes logic reasoning data designed to activate the reasoning potential of MLLMs using multimodal Chain-of-Thoughts. Sun et al.Sun et al. (2025c) propose two effective methods for synthesizing multimodal CoT datasets, offering practical solutions for generating large-scale multimodal reasoning data. For audio data, CoTAXie et al. (2025) provides multi-task audio datasets with structured Chain-of-Thought reasoning processes.
@@
Given that most current CoT datasets focus on mathematically rigorous and code-based reasoning tasks, some researchers have begun exploring CoT data for other application scenarios. This expansion is of crucial importance because it demonstrates that CoT is not merely a specialized technique applicable only to mathematical logic, but rather a general framework with wide applicability that simulates the complex cognitive processes of humans. LongPerceptualThoughtsLiao et al. (2025) aims to fill the gap in reasoning datasets for visual perception tasks, providing a dataset of 30K perceptual reasoning samples with long-form CoT traces. Meanwhile, Emma-XSun et al. (2024c) and Embodied-ReasonerZhang et al. (2025e) supply training data for embodied intelligence models. Emma-X offers a hierarchical embodied dataset containing 60K robot operation trajectories, while Embodied-Reasoner provides 9.3K coherent observation–thought–action trajectories. DriveLMM-o1 Ishaq et al. (2025) contributes over 18K visual question-answering instances for step-by-step visual reasoning in autonomous driving, covering tasks such as perception, prediction, and planning.
@@ 3.2.2 Training Paradigm
To address this limitation and enhance learning efficiency, the researchers incorporated the scaffolding theory from cognitive science and guided the model to gradually acquire the reasoning ability. For instance, LlamaV-o1Thawakar et al. (2025) adopts a multi-step curriculum learning strategy where the model first learns basic subtasks before progressing to multi-step reasoning training. Similarly, CoR(Yu et al., 2025) and LLaVA-CoT(Xu et al., 2024a) employ a Progressive Paradigm Training (PPT) approach, introducing different types of reasoning data in stages to gradually expand the model’s reasoning capabilities.
@@
Unlike conventional MLLMs that use Reinforcement Learning from Human Feedback (RLHF) to align with human preferences and reduce hallucinations, reinforcement learning in CoT-MLLMs focuses more on improving the logicality, completeness, and problem-solving effectiveness of reasoning chains. This training paradigm typically employs reinforcement fine-tuning (RFT) methods based on Direct Preference Optimization (DPO)(Rafailov et al., 2023) or Group Relative Policy Optimization (GRPO)(Shao et al., 2024b).
For example, Seg-Zero(Liu et al., 2025d) uses five manually designed reward functions to enable the model to acquire a reasonable level of reasoning ability without relying on any supervised reasoning data. Inspired by the cold-start strategy of DeepSeek-R1(Guo et al., 2025), Vision-R1(Huang et al., 2025) proposes a Progressive Thinking Suppression Training (PTST) strategy to alleviate the optimization challenge of excessive reasoning in early-stage reinforcement learning. This method restricts reasoning length initially and gradually relaxes this constraint during training, enabling the model to autonomously learn to use longer reasoning chains to solve complex problems, thereby enhancing its reasoning capabilities. In addition, Visual-RFT(Liu et al., 2025e) demonstrates that reinforcement fine-tuning can be effectively applied beyond mathematics and code to visual perception tasks, and it is the first to adopt Reinforcement Learning with Verifiable Rewards (RLVR) in image-based multimodal models. R1-Omni(Zhao et al., 2025b) further extends RLVR to video-based multimodal models for the first time.
@@
ARES(Byun et al., 2024) proposes a two-stage algorithm that alternates between reinforcement learning and supervised fine-tuning. First, a teacher model scores the outputs to guide reinforcement learning; then, the teacher corrects erroneous reasoning steps, followed by supervised fine-tuning to stabilize post-RL performance. Moreover, Insight-V(Dong et al., 2025) adopts an iterative DPO strategy in the reinforcement learning phase. By conducting multiple rounds of DPO training and sampling, the model better simulates the online DPO process, effectively addressing the issue of preference data drifting from the model distribution in offline generation. To further enhance the stability of multimodal reasoning models, Wang et al.(Wang et al., 2024e) introduce a Mixed Preference Optimization (MPO) approach, whose loss function consists of three components: preference loss, quality loss, and generation loss.